OpenStudy (anonymous):

Find the derivative. y = ln(e^(8x)) y'=

OpenStudy (barrycarter):

Surely you jest. ln(e^(8x)) is just 8x.

OpenStudy (mimi_x3):

Use: \[\frac{d}{dx} \ln(...) = \left(\frac{1}{...}\right) *(...)' \] so.. \[\frac{d}{dx} \ln(e^{8x} ) = \frac{1}{e^{8x}} *8e^{8x} =>\frac{8e^{8x}}{e^{8x}} =>8\]

OpenStudy (anonymous):

wow

OpenStudy (mimi_x3):

wow? lol

OpenStudy (anonymous):

\[y=\ln (f(x))\] \[y'=\frac{1}{f(x)}*f'(x)\]